1. The Taipei City Council has 11 members: 6 males, 5 females. They wish to form a committee of 5.

(a) (2 pts) How many ways can the city council form the committee?

(b) (12 pts) Complete the probability distribution below:

Number of Females x Probability of Number of Females P(x) 0 1 2 3 4 5 3

Number of females

x

Probability of Number of Females

P(x)

0
1
2
3
4
5

(c) (5 pts) Find the probability of having all females in the committee of the Taipei City Council. What can we conclude about the committee of the Taipei City Council of having all females? Explain why?

(d) (5 pts) Find the probability of having at least 3 females in the committee of the Taipei City Council.

(e) (7 pts) Find the probability of having at most 4 males in the committee of the Taipei City Council.

(f) (2 pts) Find μ for the number of females, round your answer to 3 decimals if possible.

(g) (2 pts) Find σ for the number of females, round your answer to 3 decimals if possible.
Answer #1

There is a very low probability of having all females.

e)P(atmost 4 males)=P(Atleast 1female)=1-P(no female)=1-0.01299=0.98701